Please refer to illustration for question.
3 -2 3 P P-AP such that A=0 زما Find a nonsingular matrix is diagonal where 0 -3 2
Answer #1

We hare 3 2 3 A = 6 3 -2 0 -3 2 3 -3 x-3 Now dit (XI-A) = 0 x-3 2. n-2 0 3 2 pr-3) 2 (7-2 (x-3) 1 x 5) 3,50 rector eigen A ar- 25-29 13, 26 - 1 2 - 2 23 2 0, T«ке Жи. - 27, + 4 + = 0, ія 5 = 5. (1 ) vector в» 1 Е. х, To be би Ру не rawe а 1 Аст го ,5 - 5 P 0 - 2 6 0 3. g dit pa - 30 -30 0 0 Adap = - 55 - 20 -2. a 30 - 55 27 0 9 -6 -2 - 2 a 30 -55 20 1 -30 0 -6 0 - 2 -2 3